Supposed that the z is the test statistics for hypothesis testing, calculate the value of the z for each of the following. Express your answer in 2 decimal places. µ = 77, σ = 6, n = 109, x̄ = 78.2
1
Expert's answer
2020-07-20T18:11:49-0400

Given:

"\overline{x}=78.2" , "\mu= 77,""\sigma= 6", n=109

formula:

"Z= \frac{\overline{x}-\mu}{\frac{\sigma}{\sqrt{n}}}"


solution:

"Z= \frac{78.2-77}{\frac{6}{\sqrt{109}}}"

=1.2/0.575

=2.087


answer: value of Z= 2.087